# Higher-order triadic percolation on random hypergraphs

### v1

@@ III.3.3 Route to chaos in higher-order triadic percolation
In Fig. 8 (c) and (d) we show two examples in which the reentrant collapsed state is observed. In panels (e) and (f) we examine the condition of having a non-trivial dynamics for corresponding cases. We reveal that the collapsed state emerge exactly when the mentioned condition is fulfilled.
Figure 10: The dependence of the order parameter $R$ of HHOTI versus $p$ is shown for hypegraphs having exclusively positive HHOTIs. S Panel (a) shows the dependence of the phase diagram with the number of hierarchical layers $L$, panel (b) shows the dependence with the value $m$ of the hyperedge cardinality. Solid lines indicate theoretical predictions symbols indicate the results of Monte Carlo simulations on hypergraphs of $N=5\times 10^{4}$ nodes. The structural hypergraphs have a Poisson hyperdegree distribution P(k) with an average $c=2$ and hyperedges of fixed cardinality $m$ ($m=8$ in panel (a)). The regulatory network is formed exclusively by positive regulations and has a Poisson regulatory degree with an average $c^{+}=2.5$ (panel (a)) and $c^{+}=5$ (panel (b)).
Figure 11: Illustrative examples of the map $h_{m,p}(R)$ representing the HHOTP dynamics for specific values of $p$ are plotted in the first and third row together with the identity line (dashed) whose intersection with the map indicates the fix point $R^{\star}=h_{m,p}(R^{\star})$. The corresponding theoretical phase diagrams/orbit diagrams of HHOTI are plotted in the second and fourth row. Panel (a) shows the HHOTP with exclusive positive regulations. Panels (b), (c), (g) and (h) show the HHOTP with exclusive negative regulations, and panel (i) shows an example of the general case with both positive and negative regulations. The different panels correspond to different phase transitions: a discontinuous transition (d, f), a continuous transition (e), a period-2 bifurcations (j) and a route to chaos (k,l). The random hypergraphs have a Poisson hyperdegree distribution with an average $c$ and a Poisson regulatory degree distribution with an average $c^{+}$ and $c^{-}$ for all layers. The hyperedges have a fixed cardinality $m$. The hierarchical regulations have $L$ layers. The parameters used in the figure are: $c=3$, $c^{+}=5$, $m=5$, $L=10$ (panel (a)), $c=5$, $c^{-}=3$, $m=10$, $L=2$ (panel (b)), $c=3$, $c^{-}=10$, $m=2$, $L=50$ (panel (c)), $c=3$, $c^{-}=2$, $m=2$, $L=4$ (panel (j)), $c=3$, $c^{-}=5$, $m=3$, $L=100$ (panel (k)), $c=30$, $c^{+}=20$, $c^{-}=30$, $m=10$, $L=10$ (panel (l)).
@@ IV Hierarchical higher-order triadic percolation
Assuming that the nested HHOTIs regulate the interactions simultaneously, we define the HHOTP by a two steps iterative dynamics, which differs from the dynamics of HOTP only by a modification of Step 2. Indeed Step 2 becomes
@@
    An active regulatory interaction in $W^{[\mu]}$ requires that it is not down-regulated by regulatory interactions in $W^{[\mu+1]}$.
@@
for $\mu\in\{2,\ldots,L\}$ with $p_{H,[L+1]}=1$. Note that here the generating function $G_{0,\ell}^{\pm}\left(x\right)$ is given by Eq. 4. These equations that implement Step 2 can be encoded into a map
@@
while the equations that implement Step 1 remain the same as for HOTP and obey Eq. 9, i.e. they are given by $R^{(t)}=f_{m}\left(p_{H}^{(t-1)}\right)$. Thus we obtain that the dynamics of HHOTP can be encoded in a one-dimensional map
@@
This map, in presence of negative regulations is no longer unimodal, thus HHOTP displays a phenomenology that is significantly different from HOTP and displays a route to chaos that is not any longer in the universality class of the logistic map. Here we discuss the rich phenomenology of this complex dynamical process by considering first the simple case in which only positive regulations are present, then the case in which only negative regulations are present and finally the general case in which both positive and negative regulations are present with non-zero probability.
When only positive regulations are present, similarly to what happens for HOTP under the same conditions, HHOTP displays a discontinuous hybrid transition. In order to illustrate this phenomenology in a simple case we consider a random hypergraph with fixed hyperedge cardinality $m$, a Poisson hyperdegree distribution $P(k)$, and Poisson regulatory degree distribution $\hat{P}_{[\mu]}(\kappa^{\pm})$ of a same average degree across all layers. In this scenario we observe that the hypergraphs with a larger hyperedge cardinality $m$ and a smaller number of regulatory layers $L$ are more robust, as indicated by their lower critical percolation threshold $p_{c}$, and their smaller discontinuity $R^{*}(p_{c})$ (see Fig. 10).
When only negative regulations are present, HHOTP displays a rich phenomenology and the nature of the dynamics changes drastically with respect to HOTP. We recall that when only negative regulations are present, HOTP displays either continuous transitions or period-2 oscillations. On contrast, HHOTP can display continuous transitions, discontinuous hybrid transitions, periodic oscillations, and chaos (see Fig. 11). Specifically, as long as there negative regulatory interactions and more than one layers, i.e. $L>1$, the map $h_{m,p}(R)$ can have multiple local maxima and minima in general, and the order parameter of HHOTP undergoes a route to chaos no longer in the universality class of the logistic map (see Fig. 11).
@@ V.1 Interdependent higher-order triadic percolation (IHOTP)
The connectivity of hypergraphs can be probed not only with hypergraph percolation resulting from the random deactivation of hyperedges, but also with several other higher-order percolation processes [14], in which the activation of a hyperedge requires, in addition to the regulatory rules, also cooperation among all nodes that belong to it.
In interdependent hyperedge percolation the interdependent hypergraph giant component is the extensive connected component whose nodes and hyperedges satisfy the following recursive self-consistent conditions.
@@
Here we define interdependent higher-order triadic percolation (IHOTP) which generalizes HOTP. As HOTP, IHOTP is defined in terms of a two step iterative process, the first step implementing percolation of the structural hypergraph and the second step implementing regulation of the hyperedges. The difference among HOTP and IHOTP is that in Step 1 of IHOTP we consider interdependent hyperedge percolation instead of simple hyperedge percolation. Specifically, in IHOTP Step 1 of the iterative algorithm is modified as:
@@
The equation determining the regulation process occurring at Step 2 is unchanged with respect to HOTP and is given by Eq. 8 that we rewrite here for convenience,
@@
Figure 12: The effect due to the discontinuous hybrid transition of interdependent hyperedge percolation on the dynamical properties of IHOTP. The figure displays the maps $g_{p}(R)$, $f_{m}(p_{H})$ determining Step 1 and Step 2 of the IHOTP dynamics (panels (a)-(c)), the one-dimensional map $h_{m,p}(R)=f_{m}(g_{p}(R))$ encoding for the combined Step 1 and Step 2 dynamics (panels (d)-(f)) and their corresponding phase diagrams/orbit diagrams (panels (g)-(i)). The paneld correspond to parameter values in which the interdependent hyperedge percolation (function $R=g_{p}(p_{H})$) displays either a continuous (panels (a)-(d)-(g)) or a discontinuous phase transition (panels (b),(c),(e),(f),(h),(i)). In panels (a)-(f) the dashed line indicate either the discontinuous jump of intedependent percolation (panels (a)-(c)) or the values of $R$ that cannot be reached by the IHOTP dynamics (panels (d)-(f)). The horizontal line in panels (a)-(c) shows the value $R=\tilde{R}$ at which the function $p_{H}=g_{p}(R)$ reaches its local maximum. The structural hypergraphs have Poisson hyperdegree distribution with average hyperdegree $c=10$ and hyperedges with cardinality distribution $Q(m)$ given by Eq. 38. Here $r$ indicates the ratio between the number of hyperedges of cardinality $m=2$ and the number of hyperedges of cardinality $m=3$. The hypergraphs have regulatory degree distributions with average degrees $c^{+}=4$ and $c^{-}=4.5$. The adopted value of $r$ are $r=1$ (panels (a),(d),(g)), $r=0.6$ (panels (b),(e),(h)) and $r=0.4$ (panels (c),(e),(i)). The critical value of $r=r_{H}^{c}$ of this model is $r_{H}^{c}=2/3$ see Eq. 40. Panels (a)-(f) are plotted for $p=0.6$.
@@
In order to provide concrete evidence of the effect due to the discontinuous hybrid transition of interdependent hyperedge percolation on the dynamical properties of IHOTP, let us again write Eq. 34 and Eq. 35 in the form of one-dimensional maps
@@
We consider structural hypergraphs with a Poisson hyperdegree distribution $P(k)$ with average hyperdegree $c$ and hyperedges of cardinality given either by $m=2$ or $m=3$. Specifically we consider the cardinality distribution $Q(m)$ given by
@@
where we denote with $r$ the ratio between the number of hyperedges of cardinality $m=2$ and the number of hyperedges of cardinality $m=3$, i.e. $r=q_{2}/q_{3}$. It is known [14] that in this ensembles of hypegraphs, interdependent hypergraph percolation displays a tricritical point at $r=r_{H}^{c}$, $p=p_{H}^{c}$ given by
@@
It is thus instructive to investigate the behavior of IHOTP in this scenario (see Fig. 12). Recall that the dynamics of IHOTP encoded in the map $h_{m,p}(R)$ defined in Eq. 37 is initialized as that at time $t=0$, when all hyperedges are active with probability $p_{H}^{0}$. When the percolation transition is discontinuous, there will be an interval of values of $R$ from 00 to a finite value $\mathcal{R}$ that is not physical, i.e. cannot be achieved (see Fig. 12 (b) and (c)). Thus the corresponding map $h_{m,p}(R)$, is not defined in the entire range $[0,1]$ as it does not have compact support (see Fig. 12 (e) and (f)). This effect significantly changes the dynamical properties of IHOTP. In particular the map $h_{m,p}(R)$ might no longer display a local maximum. As discussed in the case of HOTP, (see Appendix A) in the situation in which $f_{m}(p_{H})$ is a continuous function, i.e. when the percolation transition is continuous, the map $h_{m,p}(R)$ displays a local maximum for $R=\tilde{R}$ where $g_{p}^{\prime}(\tilde{R})=0$. Therefore the position of this maximum is independent of the percolation process under consideration. Specifically, for Poisson positive and negative regulatory degree distributions with averages $c^{+}$ and $c^{-}$ we have that $\tilde{R}$ is given by Eq. 28 as for HOTP. When the considered percolation process displays instead a discontinuous phase transition at $p=p_{c}$ with a discontinuous jump $R=\mathcal{R}$ we cannot observe a local maximum of the map $h_{m,p}(R)$ as long as
@@
The critical value of $r$ at which we have $\mathcal{R}=\tilde{R}$ is here indicated as $r=r_{c}$. Thus for $r<r_{c}$ we cannot observe a route to chaos of the order parameter of IHOTP and only stationary states and period-2 oscillations are allowed. In Fig. 16 we show that $r_{c}$ is an increasing function of the average negative regulatory degree $c^{-}$. This can be explained as follows. First we observe that $\tilde{R}$ given by (Eq. 28) is a decreasing function of $c^{-}$. This indicates that a smaller discontinuity of the interdependent percolation order parameter will be enough to impede the route to chaos. Thus this it is consistent with the monotonically increasing behavior of $r_{c}$ as a function of $c^{-}$.
Figure 13: The critical value $r_{c}$ (satisfying Eq. 40) of the fraction between the number of hyperedges of cardinality $m=2$ and the number of hyperedges of cardinality $m=3$, is plotted as a function of the average negative regulatory degree $c^{-}$. Thus if if $r<r_{c}$ a route to chaos cannot be observed in IHOTP. The plot is obtained by considering a random hypergraph with Poisson hyperdegree distribution with an average hyperdegree $c=5$, cardinality distribution $Q(m)$ given by Eq. 38 and the regulatory interactions have a Poisson distribution with an average $c^{+}=5$ and $c^{-}$.
@@ V.2 Generalization to regulated nodes
An important generalization of HOTP is higher-order node dynamic percolation (HONDP) where the nodes are regulated instead of the hyperedges. In the context of networks, the regulation of the nodes instead of the edges has been already considered in Ref. [13] where it was demonstrated that this generalization of triadic percolation still leads to a route to chaos of the order parameter. Here we explore HONDP on hypergraphs emphasizing the peculiar properties of this dynamical percolation process. First of all we observe a major difference between regulating hyperedges and regulating nodes: while hyperedges allows can be hierarchically regulated, nodes cannot. Thus we cannot define the node-regulation version of HHOTP. The second observation that we make is that HONDP admits two formulations that depend on the higher-order nature of the hypergraph and have no correspondence in HOTP. Indeed in hypergraphs, the effect of down-regulating a hyperedge is deterministic: it simply deactivates the hyperedge. Instead down-regulating a node can lead to two possible consequences. In hyperedge percolation [14] with random deactivation of the nodes, all the hyperedges involving the down-regulated nodes will reduce their cardinality by one. For instance, in social networks, a meeting can continue to take place also if one participant leaves. Alternatively, down-regulating a node might lead to deactivation of all the hyperedges the node belongs to [52]. This for instance occurs in chemical reactions where if a reactant is missing the reaction cannot take place. Here we call this process cooperative hypergraph percolation (CHP) (see Fig. 14 and Ref. [19] for further details). According to these two definitions of node hypergraph percolation, we can define two versions of HONDP.
@@
  • Step 1: Given the set of active nodes at time $t-1$, a node is considered active at time $t$ if it is not deactivated at time $(t-1)$ and if he belongs to the giant component of (a) hyperedge percolation or (b) cooperative hyperedge percolation.
@@
Figure 14: Schematic diagrams demonstrating the effect of node down-regulation in simple and in cooperative node hyperedge percolation. Panel (a) shows a hypergraph. Panel (b) shows the effect of node deactivation in simple node hyperedge percolation. When a node is removed (deactivated), the cardinality of the hyperedges it belongs to decreases. Panel (c) shows the effect of node deactivation in cooperative node hyperedge percolation. When a node is deactivated, all hyperedges it belongs to are also deactivated.
The equations determining the size of the giant component $R^{(t)}$ given the probability $p_{N}^{(t-1)}$ that the nodes are active depend on the considered percolation process. In the case (a) Step 1 implements hyperedge percolation with regulation of the nodes. In this model nodes and hyperedges satisfy the following self-consistent and recursive conditions [14].
@@
The equations for hypergraphs [14] can be used to obtain the size of the giant component $R^{(t)}$ at of HONDP at time $t$ given the probability $p_{N}^{(t-1)}$ that nodes are not down-regulated at time $t-1$. These equations are given by:
@@
      at least one of its node are in the giant component.
@@
In both cases Step 2 is enforced by the equation implementing the regulation of the nodes, and thus expressing the probability $p_{N}^{(t)}$ that a node is active as a function of the probability $R^{(t)}$ that its regulator nodes are in the giant component at Step 1, which is given by,
@@
The dynamics of HONDP, as the dynamics of HOTP, is described by a percolation order parameter $R=R^{(t)}$ that, in presence of both positive and negative regulatory interactions undergoes a route to chaos in the universality class of the logistic map, in both versions of the model. To show this, let us adopt the same notation as for HOTP, and express the results of Step 1 given by the Eqs. 42 in case (a) and Eqs. 43 in case (b) as
@@
while we express the Eq. 44 implementing Step 2 as
@@
With this notation, the dynamics of HONDP can be reduced to the one dimensional map
@@
As for HOTP, also for HONDP this map is continuous and since the function $f_{m}(R)$ is monotonous, the map $h_{m,p}(R)$ displays a maximum only when the function $p_{N}=g_{p}(R)$ displays a maximum. This implies that only if both negative and positive regulatory interactions are present with finite probability, HONDP displays a route to chaos of its order parameter.
@@
In order to illustrate some specific properties of HONDP we consider here the limiting case in which there are only positive regulations. In this scenario, similary to what we have previously discussed for HOTP, HONDP displays exclusively discontinuous hybrid transitions.
@@
We first compare case (i) of HONP with HOTP. In both cases Step 1, implements hypergraph percolation, however in HONDP the nodes are upregulated with probability $p_{N}$, while in HOTP the hyperedges are upregulated with probability $p_{H}$. Interestingly, when regulatory interactions are not present, i.e. $p_{N}=p_{H}=p$, hypergraph percolation with deactivation of nodes and hypergraph percolation with deactivation of hyperedges displays a continuous phase transition at the same critical threshold [14]. However, in the presence of exclusive positive regulatory interactions, HONDP (i) and HOTP display discontinuous transitions and HONDP (i) has a larger critical threshold than HOTP, indicating a less robust giant component (see Fig. 15).

### v2

@@ III.3.5 Reentrant orbit diagram in HOTP
In Fig. 9 (c) and (d) we show two examples in which the reentrant collapsed state is observed. In panels (e) and (f) we examine the condition of having a non-trivial dynamics for corresponding cases. We reveal that the collapsed state emerges exactly when the mentioned condition is fulfilled.
Figure 11: The dependence of the order parameter $R$ of HHOTI versus $p$ is shown for hypergraphs having exclusively positive HHOTIs. Panel (a) shows the dependence of the phase diagram with the number of hierarchical layers $L$, and panel (b) shows the dependence with the value $m$ of the hyperedge cardinality. Solid lines indicate theoretical predictions symbols indicate the results of Monte Carlo simulations on hypergraphs of $N=5\times 10^{4}$ nodes. The structural hypergraphs have a Poisson hyperdegree distribution P(k) with an average $c=2$ and hyperedges of fixed cardinality $m$ ($m=8$ in panel (a)). The regulatory network is formed exclusively by positive regulations and has a Poisson regulatory degree with an average $c^{+}=2.5$ (panel (a)) and $c^{+}=5$ (panel (b)).
Figure 12: Illustrative examples of the map $h_{m,p}(R)$ representing the HHOTP dynamics for specific values of $p$ are plotted in the first and third row together with the identity line (dashed) whose intersection with the map indicates the fixed point $R^{\star}=h_{m,p}(R^{\star})$. The corresponding theoretical phase diagrams/orbit diagrams of HHOTI are plotted in the second and fourth rows. Panel (a) shows the HHOTP with exclusive positive regulations. Panels (b), (c), (g), and (h) show the HHOTP with exclusive negative regulations, and panel (i) shows an example of the general case with both positive and negative regulations. The different panels correspond to different phase transitions: a discontinuous transition (d, f), a continuous transition (e), a period-2 bifurcation (j), and a route to chaos (k,l). The random hypergraphs have a Poisson hyperdegree distribution with an average $c$ and a Poisson regulatory degree distribution with an average $c^{+}$ and $c^{-}$ for all layers. The hyperedges have a fixed cardinality $m$. The hierarchical regulations have $L$ layers. The parameters used in the figure are: $c=3$, $c^{+}=5$, $m=5$, $L=10$ (panel (a)), $c=5$, $c^{-}=3$, $m=10$, $L=2$ (panel (b)), $c=3$, $c^{-}=10$, $m=2$, $L=50$ (panel (c)), $c=3$, $c^{-}=2$, $m=2$, $L=4$ (panel (j)), $c=3$, $c^{-}=5$, $m=3$, $L=100$ (panel (k)), $c=30$, $c^{+}=20$, $c^{-}=30$, $m=10$, $L=10$ (panel (l)).
@@ IV Hierarchical higher-order triadic percolation
Assuming that the nested HHOTIs regulate the interactions simultaneously, we define the HHOTP by two-step iterative dynamics, which differs from the dynamics of HOTP only by a modification of Step 2. Indeed Step 2 becomes
@@
    An active regulatory interaction in $W^{[\mu]}$ requires that it is not regulated by any active negative regulator via active regulatory interactions in $W^{[\mu+1]}$ and is regulated by at least one active positive regulator via active regulatory interactions in $W^{[\mu+1]}$.
@@
for $\mu\in\{2,\ldots,L\}$ with $p_{H,[L+1]}=1$. Note that here the generating function $G_{0,\ell}^{\pm}\left(x\right)$ is given by Eq. (4). These equations that implement Step 2 can be encoded into a map
@@
while the equations that implement Step 1 remain the same as for HOTP and obey Eq. (9), i.e. they are given by $R^{(t)}=f_{m}\left(p_{H}^{(t-1)}\right)$. Thus we obtain that the dynamics of HHOTP can be encoded in a one-dimensional map
@@
This map in the presence of negative regulations is no longer unimodal, thus HHOTP displays a phenomenology that is significantly different from HOTP and displays a route to chaos that is no longer in the universality class of the logistic map. Here we discuss the rich phenomenology of this complex dynamical process by considering first the simple case in which only positive regulations are present, then the case in which only negative regulations are present and finally the general case in which both positive and negative regulations are present with non-zero probability.
When only positive regulations are present, similarly to what happens for HOTP under the same conditions, HHOTP displays a discontinuous hybrid transition. In order to illustrate this phenomenology in a simple case we consider a random hypergraph with fixed hyperedge cardinality $m$, a Poisson hyperdegree distribution $P(k)$, and Poisson regulatory degree distribution $\hat{P}_{[\mu]}(\kappa^{\pm})$ of a same average degree across all layers. In this scenario, we observe that the hypergraphs with a larger hyperedge cardinality $m$ and a smaller number of regulatory layers $L$ are more robust, as indicated by their lower critical percolation threshold $p_{c}$, and their smaller discontinuity $R^{*}(p_{c})$ (see Fig. 11).
When only negative regulations are present, HHOTP displays a rich phenomenology and the nature of the dynamics changes drastically with respect to HOTP. We recall that when only negative regulations are present, HOTP displays either continuous transitions or period-2 oscillations. In contrast, HHOTP can display continuous transitions, discontinuous hybrid transitions, periodic oscillations, and chaos (see Fig. 12). Specifically, as long as there are negative regulatory interactions and more than one layer, i.e. $L>1$, the map $h_{m,p}(R)$ can have multiple local maxima and minima in general, and the order parameter of HHOTP undergoes a route to chaos no longer in the universality class of the logistic map (see Fig. 12).
@@ V.1 Interdependent higher-order triadic percolation (IHOTP)
The connectivity of hypergraphs can be probed not only with hypergraph percolation resulting from the random deactivation of hyperedges but also with several other higher-order percolation processes [14], in which the activation of a hyperedge requires, in addition to the regulatory rules, also cooperation among all nodes that belong to it.
In interdependent hyperedge percolation, the interdependent hypergraph giant component is the extensive connected component whose nodes and hyperedges satisfy the following recursive self-consistent conditions.
@@
Here we define interdependent higher-order triadic percolation (IHOTP) which generalizes HOTP. As HOTP, IHOTP is defined in terms of a two-step iterative process, the first step implementing percolation of the structural hypergraph and the second step implementing regulation of the hyperedges. The difference between HOTP and IHOTP is that in Step 1 of IHOTP, we consider interdependent hyperedge percolation instead of simple hyperedge percolation. Specifically, in IHOTP Step 1 of the iterative algorithm is modified as:
@@
The equation determining the regulation process occurring at Step 2 is unchanged with respect to HOTP and is given by Eq. (8) that we rewrite here for convenience,
@@
Figure 13: The effect due to the discontinuous hybrid transition of interdependent hyperedge percolation on the dynamical properties of IHOTP. The figure displays the maps $g_{p}(R)$, $f_{m}(p_{H})$ determining Step 1 and Step 2 of the IHOTP dynamics (panels (a)-(c)), the one-dimensional map $h_{m,p}(R)=f_{m}(g_{p}(R))$ encoding for the combined Step 1 and Step 2 dynamics (panels (d)-(f)) and their corresponding phase diagrams/orbit diagrams (panels (g)-(i)). The panel corresponds to parameter values in which the interdependent hyperedge percolation (function $R=g_{p}(p_{H})$) displays either a continuous (panels (a)-(d)-(g)) or a discontinuous phase transition (panels (b),(c),(e),(f),(h),(i)). In panels (a)-(f) the dashed lines indicate either the discontinuous jump of interdependent percolation (panels (a)-(c)) or the values of $R$ that cannot be reached by the IHOTP dynamics (panels (d)-(f)). The horizontal line in panels (a)-(c) shows the value $R=\tilde{R}$ at which the function $p_{H}=g_{p}(R)$ reaches its local maximum. The structural hypergraphs have Poisson hyperdegree distribution with average hyperdegree $c=10$ and hyperedges with cardinality distribution $Q(m)$ given by Eq. (38). Here $r$ indicates the ratio between the number of hyperedges of cardinality $m=2$ and the number of hyperedges of cardinality $m=3$. The hypergraphs have regulatory degree distributions with average degrees $c^{+}=4$ and $c^{-}=4.5$. The adopted value of $r$ are $r=1$ (panels (a),(d),(g)), $r=0.6$ (panels (b),(e),(h)) and $r=0.4$ (panels (c),(e),(i)). The critical value of $r=r_{H}^{c}$ of this model is $r_{H}^{c}=2/3$ see Eq. (40). Panels (a)-(f) are plotted for $p=0.6$.
@@
In order to provide concrete evidence of the effect due to the discontinuous hybrid transition of interdependent hyperedge percolation on the dynamical properties of IHOTP, let us again write Eq. (34) and Eq. (35) in the form of one-dimensional maps
@@
We consider structural hypergraphs with a Poisson hyperdegree distribution $P(k)$ with average hyperdegree $c$ and hyperedges of cardinality given either by $m=2$ or $m=3$. Specifically, we consider the cardinality distribution $Q(m)$ given by
@@
where we denote with $r$ the ratio between the number of hyperedges of cardinality $m=2$ and the number of hyperedges of cardinality $m=3$, i.e. $r=q_{2}/q_{3}$. It is known [14] that in this ensemble of hypergraphs, interdependent hypergraph percolation displays a tricritical point at $r=r_{H}^{c}$, $p=p_{H}^{c}$ given by
@@
It is thus instructive to investigate the behavior of IHOTP in this scenario (see Fig. 13). Recall that the dynamics of IHOTP encoded in the map $h_{m,p}(R)$ defined in Eq. (37) is initialized as that at time $t=0$, when all hyperedges are active with probability $p_{H}^{0}$. When the percolation transition is discontinuous, there will be an interval of values of $R$ from 00 to a finite value $\mathcal{R}$ that is not physical, i.e. cannot be achieved (see Fig. 13 (b) and (c)). Thus the corresponding map $h_{m,p}(R)$, is not defined in the entire range $[0,1]$ as it does not have compact support (see Fig. 13 (e) and (f)). This effect significantly changes the dynamical properties of IHOTP. In particular the map $h_{m,p}(R)$ might no longer display a local maximum. As discussed in the case of HOTP, (see Appendix A) in the situation in which $f_{m}(p_{H})$ is a continuous function, i.e. when the percolation transition is continuous, the map $h_{m,p}(R)$ displays a local maximum for $R=\tilde{R}$ where $g_{p}^{\prime}(\tilde{R})=0$. Therefore the position of this maximum is independent of the percolation process under consideration. Specifically, for Poisson positive and negative regulatory degree distributions with averages $c^{+}$ and $c^{-}$ we have that $\tilde{R}$ is given by Eq. (28) as for HOTP. When the considered percolation process displays instead a discontinuous phase transition at $p=p_{c}$ with a discontinuous jump $R=\mathcal{R}$ we cannot observe a local maximum of the map $h_{m,p}(R)$ as long as
@@
The critical value of $r$ at which we have $\mathcal{R}=\tilde{R}$ is here indicated as $r=r_{c}$. Thus for $r<r_{c}$ we cannot observe a route to chaos of the order parameter of IHOTP and only stationary states and period-2 oscillations are allowed. In Fig. 14 we show that $r_{c}$ is an increasing function of the average negative regulatory degree $c^{-}$. This can be explained as follows. First, we observe that $\tilde{R}$ given by (Eq. (28)) is a decreasing function of $c^{-}$. This indicates that a smaller discontinuity of the interdependent percolation order parameter will be enough to impede the route to chaos. Thus this is consistent with the monotonically increasing behavior of $r_{c}$ as a function of $c^{-}$.
Figure 14: The critical value $r_{c}$ (satisfying Eq. (40)) of the fraction between the number of hyperedges of cardinality $m=2$ and the number of hyperedges of cardinality $m=3$, is plotted as a function of the average negative regulatory degree $c^{-}$. Thus if $r<r_{c}$ a route to chaos cannot be observed in IHOTP. The plot is obtained by considering a random hypergraph with Poisson hyperdegree distribution with an average hyperdegree $c=5$, cardinality distribution $Q(m)$ given by Eq. (38) and the regulatory interactions have a Poisson distribution with an average $c^{+}=5$ and $c^{-}$.
@@ V.2 Generalization to regulated nodes
An important generalization of HOTP is higher-order node dynamic percolation (HONDP) where the nodes are regulated instead of the hyperedges. The higher-order nature of this process is due to the fact that each node is typically regulated by several regulator nodes. In the context of networks, the regulation of the nodes instead of the edges has been already considered in Ref. [13] where it was demonstrated that this generalization of triadic percolation still leads to a route to chaos of the order parameter. Here we explore HONDP on hypergraphs emphasizing the peculiar properties of this dynamical percolation process and its relation to HOTP. First of all, we observe a major difference between regulating hyperedges and regulating nodes: while hyperedges can be hierarchically regulated, nodes cannot. Thus we cannot define the node-regulation version of HHOTP. The second observation that we make is that HONDP admits two formulations that depend on the higher-order nature of the hypergraph and have no correspondence in HOTP. Indeed in hypergraphs, the effect of down-regulating a hyperedge is deterministic: it simply deactivates the hyperedge. Instead down-regulating a node can lead to two possible consequences. In hyperedge percolation [14] with random deactivation of the nodes, all the hyperedges involving the down-regulated nodes will reduce their cardinality by one. For instance, in social networks, a meeting can continue to take place also if one participant leaves. Alternatively, down-regulating a node might lead to the deactivation of all the hyperedges the node belongs to [19]. This for instance occurs in chemical reactions where if a reactant is missing the reaction cannot take place. Here we call this process cooperative hypergraph percolation (CHP) (see Fig. 15 and Ref. [19] for further details). According to these two definitions of node hypergraph percolation, we can define two versions of HONDP.
@@
  • Step 1: Given the set of active nodes at time $t-1$, a node is considered active at time $t$ if it is not deactivated at time $(t-1)$ and if it belongs to the giant component of (a) hyperedge percolation or (b) cooperative hyperedge percolation.
@@
Figure 15: Schematic diagrams demonstrating the effect of node down-regulation in simple and cooperative node hyperedge percolation. Panel (a) shows a hypergraph. Panel (b) shows the effect of node deactivation in simple node hyperedge percolation. When a node is removed (deactivated), the cardinality of the hyperedges it belongs to decreases. Panel (c) shows the effect of node deactivation in cooperative node hyperedge percolation. When a node is deactivated, all hyperedges it belongs to are also deactivated.
The equations determining the size of the giant component $R^{(t)}$ given the probability $p_{N}^{(t-1)}$ that the nodes are active depend on the considered percolation process. In the case (a), Step 1 implements hyperedge percolation with the regulation on the nodes. In this model nodes and hyperedges satisfy the following self-consistent and recursive conditions [14].
@@
The equations for hypergraphs [14] can be used to obtain the size of the giant component $R^{(t)}$ of HONDP at time $t$ given the probability $p_{N}^{(t-1)}$ that nodes are not down-regulated at time $t-1$. These equations are given by:
@@
      at least one of its nodes is in the giant component.
@@
In both cases, Step 2 is enforced by the equation implementing the regulation of the nodes, and thus expressing the probability $p_{N}^{(t)}$ that a node is active as a function of the probability $R^{(t)}$ that its regulator nodes are in the giant component at Step 1, which is given by,
@@
The dynamics of HONDP, as the dynamics of HOTP, is described by a percolation order parameter $R=R^{(t)}$ that, in the presence of both positive and negative regulatory interactions undergoes a route to chaos in the universality class of the logistic map, in both versions of the model. To show this, let us adopt the same notation as for HOTP, and express the results of Step 1 given by the Eqs. (42) in case (a) and Eqs. (43) in case (b) as
@@
while we express the Eq. (44) implementing Step 2 as
@@
With this notation, the dynamics of HONDP can be reduced to the one-dimensional map
@@
As for HOTP, also for HONDP, this map is continuous and since the function $f_{m}(R)$ is monotonous, the map $h_{m,p}(R)$ displays a maximum only when the function $p_{N}=g_{p}(R)$ displays a maximum. This implies that only if both negative and positive regulatory interactions are present with finite probability, HONDP displays a route to chaos of its order parameter.
@@
In order to illustrate some specific properties of HONDP we consider here the limiting case in which there are only positive regulations. In this scenario, similar to what we have previously discussed for HOTP, HONDP displays exclusively discontinuous hybrid transitions.
@@
We first compare case (i) of HONP with HOTP. In both cases Step 1, implements hypergraph percolation, however, in HONDP the nodes are upregulated with probability $p_{N}$, while in HOTP the hyperedges are upregulated with probability $p_{H}$. Interestingly, when regulatory interactions are not present, i.e. $p_{N}=p_{H}=p$, hypergraph percolation with the deactivation of nodes and hypergraph percolation with the deactivation of hyperedges displays a continuous phase transition at the same critical threshold [14]. However, in the presence of exclusive positive regulatory interactions, HONDP (i) and HOTP display discontinuous transitions and HONDP (i) has a larger critical threshold than HOTP, indicating a less robust giant component (see Fig. 16).